A charged paint is spread in a very thin uniform layer over the surface of a plastic sphere of diameter 18.0 cm , giving it a charge of-29.0 μC .

A Find the electric field just inside the paint layer.

Express your answer with the appropriate units. Enter positive value if the field is directed radially inward and negative value if the field is directed radially outward.

B   Find the electric field just outside the paint layer.

Express your answer with the appropriate units. Enter positive value if the field is directed radially inward and negative value if the field is directed radially outward

C Find the electric field 6.00

cm outside the surface of the paint layer.

Express your answer with the appropriate units. Enter positive value if the field is directed radially inward and negative value if the field is directed radially outward.
Answer #1

A) Einside = 0 (since there is no charge inside the sphere)

B) E_surface = k*Q/R^2

= 9*10^9*29*10^-6/0.09^2

= 3.22*10^7 N/c (inward)

C) E_outside = k*Q/(R + h)^2

= 9*10^9*29*10^-6/(0.09 + 0.06)^2

= 1.16*10^7 N/c (inward)